What is the resistance of gold wire at 150°C , if it is 25.0 m long with a 0.100 mm diameter and has a resistance of 77.7 Ω at 20.0°C ?
Answer #1

Length of gold wire, L-25 m, diameter, d =0.1×10-3 m Resistance, R 77.7 temperature, 20°c 293K,T, -150°c-423K PL Resistance,RA Resistivity at 20°C, ρ 77.7x 3.14x(5x10-5 2.44x10- 1 =2.44 x10%(1+3.4x10-3(150°-20°)) 25 3.52x10- L Resistance, R , 3.52R 77.7x(1+3.4x10-3 (150-20 -112.04